1. A box has 3 netball and 6 volleyball cards. What is the probability of selecting...

( a) a netball card , keeping it out , and then selecting another netball card ?

( b) a netball card, keeping it out, and then selecting a volleyball card?


1
Expert's answer
2021-06-25T08:41:03-0400

(a) Probability of a selecting a netball card, keeping it out, and then selecting another netball card.

Step 1: Probability of selecting a netball card and keeping it is 3/9
Step 2: Probability of selecting another netball card 2/8
Step 3: Probability of selecting a netball card, keeping it and selecting another netball card is 
              3/9*2/8 = 6/72 = 1/12
               Answer: 1/12
              


(b) Probability of selecting a netball card, keeping it out, and then selecting a volleyball card.

Step 1: Probability of selecting a netball card and keeping it is 3/9
Step 2: Probability of selecting a volleyball card is 6/8
Step 3: Probability of selecting a netball card, keeping it and then selecting a volleyball card is 
             3/9* 6/8 = 18/72= 1/4
             Answer: 1/4